# Advances in the Shannon Capacity of Graphs

### v2

@@ 5.3 Third approach
Path graphs are universal, and for an even $\ell\geq 2$, the capacities of $\operatorname{\mathsf{P}}_{\ell-1}$ and $\operatorname{\mathsf{P}}_{\ell}$ coincide (see Theorem 2.24). The latter holds because path graphs are bipartite and therefore perfect, so their Shannon capacities coincide with their independence numbers, and also $\alpha(\operatorname{\mathsf{P}}_{\ell-1})=\frac{\ell}{2}=\alpha(\operatorname{\mathsf{P}}_{\ell})$ if $\ell\geq 2$ is even. by universality of path graphs and since the independence number of $\operatorname{\mathsf{P}}_{\ell}$ and $\operatorname{\mathsf{P}}_{\ell-1}$ coincide for an even $\ell\geq 2$ (see Theorem 2.24), for every $k\in\mathbb{N}$,
$\alpha((\mathsf{H}+\operatorname{\mathsf{P}}_{\ell-1})^{k})$ $=\sum_{i=0}^{k}\binom{k}{i}\alpha(\mathsf{H}^{i}\boxtimes\operatorname{\mathsf{P}}_{\ell-1}^{k-i})$
@@
Consequently, by raising both sides of (153) to the power $\frac{1}{k}$, and letting $k\to\infty$, it follows that $\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell-1})=\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})$. Combining the last equality with (152), hence gives
@@
Thus, by combining (153) and (155), it follows that for every $k\in\mathbb{N}$
@@
$\Theta(\mathsf{G})^{k}=\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})^{k}>\alpha\bigl((\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})^{k}\bigr)=\alpha(\mathsf{G}^{k}),$ (157)
@@
This follows directly from Theorem 5.3 and Definition 4.1, by selecting $\mathsf{H}=\operatorname{\mathsf{C}}_{k}$ for an odd $k\geq 5$. In the latter case, $\alpha(\operatorname{\mathsf{C}}_{k})<\Theta{\operatorname{\mathsf{C}}_{k}}$ as required by Theorem 5.3. ∎
@@ 6 The Shannon capacity of $q$-Kneser graphs
In this section we show the exact value of the Shannon capacity of a new family of graphs, containing the Kneser graphs.
@@
Let $n,k\in\mathbb{N}$ with $k\leq n$, let $p$ be a prime, let $q=p^{m}$ with $m\in\mathbb{N}$, and let $\mathbb{F}_{q}$ be the Galois field of order $q$. The Gaussian coefficient, denoted by $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$, is given by
@@
The Gaussian coefficient can be given the following interpretation. Let $V$ be an $n$-dimensional vector space over $\mathbb{F}_{q}$. The number of $k$-dimensional distinct subspaces that $V$ possesses is equal to $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$. We therefore define $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046pt0}\\ {\hskip-8.5359pt0\hskip-14.22636pt}\end{array}\Bigr]_{q}\triangleq 1$.
Passing to a continuous variable $q$ and letting $q$ tend to $1$ gives
@@
Let $V(n,q)$ be the $n$-dimensional vector space over the finite field $\mathbb{F}_{q}$, where $q$ is a prime power. The $q$-Kneser graph $\operatorname{\mathsf{K}}_{q}(n,k)$ is defined as follows:
  • The vertices of $\operatorname{\mathsf{K}}_{q}(n,k)$ are defined as the $k$-dimensional subspaces of $V(n,q)$.
  • Two vertices are adjacent if their intersection contains only the zero vector.
Some of the properties of the $q$-Kneser graphs are next presented (see [52, Proposition 3.4]).
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph with $n\geq 2k$ and $q$ a prime power. Then
  • $\operatorname{\mathsf{K}}_{q}(n,k)$ is of order
@@
  • $\operatorname{\mathsf{K}}_{q}(n,k)$ is of size
    $|\operatorname{\mathsf{E}}(\operatorname{\mathsf{K}}_{q}(n,k))|=\frac{1}{2}q^{k^{2}}\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-k}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}.$ (172)
  • $\operatorname{\mathsf{K}}_{q}(n,k)$ is $q^{k^{2}}\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-k}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$-regular.
  • $\operatorname{\mathsf{K}}_{q}(n,k)$ is vertex-transitive.
  • $\operatorname{\mathsf{K}}_{q}(n,k)$ is edge-transitive.
We next calculate the Shannon capacity of $\operatorname{\mathsf{K}}_{q}(n,k)$ in an analogous way to the calculation of the Shannon capacity of the Kneser graph $\operatorname{\mathsf{K}}(n,k)$ by Lovász (see [11, Theorem 13]). To that end, we derive the independence number and Lovász $\vartheta$-function of a $q$-Kneser graph, and show that they coincide, thus leading to the determination of the capacity of the graph.
@@
[29] Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph with $n\geq 2k$, and let $q$ be a prime power. Then,
@@
By the version of the Erdös-Ko-Rado Theorem for finite vector spaces (see [29, Theorem 9.8.1]), we get
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph with $n\geq 2k$ and $q$ a prime power. Then
@@
By Theorem 2.20 and since the $q$-Kneser graphs are edge-transitive (see Theorem 6.1),
@@
where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalues of the adjacency matrix of $\operatorname{\mathsf{K}}_{q}(n,k)$. By [30, Theorem 2], these eigenvalues are given by
@@
Due to the coincidence of the independence number and the Lovász $\vartheta$-function in Lemmata 6.1 and 6.2, their joint value is also equal to the Shannon capacity of the graph. This gives the following closed-form expression for the Shannon capacity of $q$-Kneser graphs.
@@
The Shannon capacity of the $q$-Kneser graph, $\operatorname{\mathsf{K}}_{q}(n,k)$, with $n\geq 2k$ and $q$ a prime power, is given by
@@ 7 A new inequality for the capacity of graphs
The following result provides a relation between the Shannon capacity of any strong product of graphs, and the capacity of the disjoint union of the component graphs. If these component graphs are connected, then their strong product is a connected graph on a number of vertices that is equal to the product of the number of vertices in each component graph, whereas the disjoint union of these component graphs is a disconnected graph on a number of vertices that is equal to the sum of the orders of the component graphs (the latter order is typically much smaller than the former). The motivation of our inequality is due to the following result:
@@
We next provide and prove the main result of this section.
@@
Let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{\ell}$ be simple graphs, then
@@
In particular, if for every $i\in[\ell]$, one of the following statements hold:
@@
thus, by combining (239) and (240), it follows that
@@
Raising both sides of the inequality to the power of $\frac{1}{k}$, and letting $k$ tend to infinity gives
@@
An anonymous reviewer brought to our attention that inequality (237) can be related to equation (1) in the preprint [47], for the complementary graph entropy, via the results of [48, Lemma 1] and [6, Equation (11.2)].
@@
with equality in (245) if the graph $\mathsf{G}$ is either self-complementary and vertex-transitive, or self-complementary and strongly regular, or either a conference graph, a Latin square graph or their complements.
@@
Next, choosing the ”diagonal” vertices of $\mathsf{G}\boxtimes\overline{\mathsf{G}}$ gives an independent set of size $n$ in $\mathsf{G}\boxtimes\overline{\mathsf{G}}$. Thus,
@@
Using (246) and (247) gives (245). Furthermore, let $\mathsf{G}$ be either (1) self-complementary and vertex-transitive or (2) strongly regular or (3) a conference graph, (4) a Latin square graph or (5) any one of the complements of the graphs in (1)–(4). Then, by Theorem 7.2 and [39, Theorems 3.23, 3.26 and 3.28], inequalities (246) and (247) hold with equality, thus inequality (245) holds with equality. ∎
@@
Inequality (245), together with a subset of the sufficient conditions for its equality in Corollary 7.1, were proved by N. Alon in [7, Theorem 2.1]. Our proof is more simple, relying on a different approach, and the original proof in [7] is also presented in Appendix E. In [7, Theorem 2.1], Alon showed that inequality (245) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive; some additional alternative sufficient conditions are provided in Corollary 7.1.
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, and let $m_{1},\ldots,m_{\ell}\in\mathbb{N}$, and let $m=\sqrt[\ell]{m_{1}\ldots m_{\ell}}$, then
@@
By Corollary 7.2 (note that the strong product in the next line involves $k$ graphs)
@@
Calculation of $m$ in the righthand side of (257) gives, by (253) and (258)–(261),
@@
This is true because if we choose $k$ that is not a multiple of $\mathrm{lcm}(q_{1},\ldots,q_{\ell})$ then there exists $i\in[\ell]$ such that $\alpha_{i}\notin\mathbb{N}$.
@@
$\vartheta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}+\frac{(\vartheta(\mathsf{G})-\sqrt{n})^{2}}{\vartheta(\mathsf{G})}.$ (265)
and for every $d$-regular graph with spectrum $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$,
$2+\frac{n-d-1}{1+\lambda_{2}}-\frac{d}{\lambda_{n}}\leq\vartheta(\mathsf{G}+\overline{\mathsf{G}})\leq\frac{n(1+\lambda_{2})}{n-d+\lambda_{2}}-\frac{n\lambda_{n}}{d-\lambda_{n}}.$ (266)
Furthermore, the following holds:
@@
    If $\mathsf{G}$ is vertex-transitive or strongly regular, then inequality (265) holds with equality.
@@
    If both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then the right inequality in (266) holds with equality.
@@
    If $\mathsf{G}$ is strongly regular, then the left inequality in (266) holds with equality.
@@
$\leq\frac{n(1+\lambda_{2})}{n-d+\lambda_{2}}-\frac{n\lambda_{n}}{d-\lambda_{n}},$ (268)
and,
@@
which proves (266). Finally, by [36, Proposition 1], if both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then the right inequality in (7) holds with equality, and if $\mathsf{G}$ is strongly regular, then the left inequality in (7) holds with equality. ∎

### v3

@@ 5.3 Third approach
Path graphs are universal, and for an even $\ell\geq 2$, the capacities of $\operatorname{\mathsf{P}}_{\ell-1}$ and $\operatorname{\mathsf{P}}_{\ell}$ coincide (see Theorem 2.24). The latter holds because path graphs are bipartite and therefore perfect, so their Shannon capacities coincide with their independence numbers, and also $\alpha(\operatorname{\mathsf{P}}_{\ell-1})=\frac{\ell}{2}=\alpha(\operatorname{\mathsf{P}}_{\ell})$ if $\ell\geq 2$ is even. By the universality of path graphs, and since the independence numbers of $\operatorname{\mathsf{P}}_{\ell}$ and $\operatorname{\mathsf{P}}_{\ell-1}$ coincide for an even $\ell\geq 2$ (see Theorem 2.24), for every $k\in\mathbb{N}$,
$\alpha((\mathsf{H}+\operatorname{\mathsf{P}}_{\ell-1})^{k})$ $=\sum_{i=0}^{k}\,\binom{k}{i}\,\alpha\bigl(\mathsf{H}^{i}\boxtimes\operatorname{\mathsf{P}}_{\ell-1}^{k-i}\bigr)$
@@
Consequently, by raising both sides of (153) to the power $\frac{1}{k}$ and letting $k\to\infty$, it follows that $\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell-1})=\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})$. Combining the last equality with (152), hence gives
@@
Thus, by combining (153) and (155), it follows that for every $k\in\mathbb{N}$,
@@
$\Theta(\mathsf{G})^{\,k}=\bigl(\Theta(\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})\bigr)^{k}>\alpha\bigl((\mathsf{H}+\operatorname{\mathsf{P}}_{\ell})^{k}\bigr)=\alpha(\mathsf{G}^{k}),$ (157)
@@
This follows directly from Theorem 5.3 and Definition 4.1, by selecting $\mathsf{H}=\operatorname{\mathsf{C}}_{k}$ for an odd $k\geq 5$. In the latter case, $\alpha(\operatorname{\mathsf{C}}_{k})<\Theta(\operatorname{\mathsf{C}}_{k})$, as required by Theorem 5.3. ∎
@@ 6 The Shannon capacity of $q$-Kneser graphs
In this section, we determine the exact Shannon capacity of the family of $q$-Kneser graphs. For the classical Kneser graphs, the Shannon capacity was determined by Lovász [11, Theorem 13] (see (53)).
@@
Let $n,k\in\mathbb{N}$ with $k\leq n$, let $p$ be a prime, let $q=p^{m}$ be a prime power with $m\in\mathbb{N}$, and let $\mathbb{F}_{q}$ denote the Galois field of order $q$. The Gaussian coefficient, denoted by $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$, is given by
@@
The Gaussian coefficient admits the following combinatorial interpretation. Let $V$ be an $n$-dimensional vector space over $\mathbb{F}_{q}$. Then, the number of distinct $k$-dimensional subspaces of $V$ is equal to $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$. Consequently, we define $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046pt0}\\ {\hskip-8.5359pt0\hskip-14.22636pt}\end{array}\Bigr]_{q}\triangleq 1$.
Letting $q$ vary continuously and taking the limit $q\to 1$ gives
@@
Let $V(n,q)$ denote the $n$-dimensional vector space over the finite field $\mathbb{F}_{q}$, where $q$ is a prime power. The $q$-Kneser graph $\operatorname{\mathsf{K}}_{q}(n,k)$ has as its vertices the $k$-dimensional subspaces of $V(n,q)$, and any two vertices are adjacent if and only if their intersection is the zero vector.
Some of the properties of the $q$-Kneser graphs are presented next (see [52, Proposition 3.4]).
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph, where $n\geq 2k$ and $q$ is a prime power. Then, the following hold:
  1. 1.
    The order of $\operatorname{\mathsf{K}}_{q}(n,k)$ is given by
@@
  2. 2.
    The size of $\operatorname{\mathsf{K}}_{q}(n,k)$ is given by
    $|\operatorname{\mathsf{E}}(\operatorname{\mathsf{K}}_{q}(n,k))|=\tfrac{1}{2}\,q^{k^{2}}\,\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-k}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}\,\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}.$ (172)
  3. 3.
    $\operatorname{\mathsf{K}}_{q}(n,k)$ is $d$-regular with $d=q^{k^{2}}\,\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-k}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}$.
  4. 4.
    $\operatorname{\mathsf{K}}_{q}(n,k)$ is vertex-transitive and edge-transitive.
Next, we determine the Shannon capacity of $\operatorname{\mathsf{K}}_{q}(n,k)$ in a manner analogous to the Lovász calculation of the Shannon capacity of the Kneser graph $\operatorname{\mathsf{K}}(n,k)$ (see [11, Theorem 13]). To that end, we derive the independence number and the Lovász $\vartheta$-function of the $q$-Kneser graph and show that they coincide, thereby yielding the Shannon capacity of the graph.
@@
[29] Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph, where $n\geq 2k$ and $q$ is a prime power. Then,
@@
By the version of the Erdős-Ko-Rado theorem for finite vector spaces (see [29, Theorem 9.8.1]), it follows that
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}_{q}(n,k)$ be a $q$-Kneser graph, where $n\geq 2k$ and $q$ is a prime power. Then,
@@
By Theorem 2.20, and since the $q$-Kneser graphs are edge-transitive (see Theorem 6.1), it follows that
@@
where $\lambda_{\max}$ and $\lambda_{\min}$ are, respectively, the largest and smallest eigenvalues of the adjacency matrix of $\operatorname{\mathsf{K}}_{q}(n,k)$. By [30, Theorem 2], these eigenvalues are given by
@@
Due to the coincidence of the independence number and the Lovász $\vartheta$-function in Lemmata 6.1 and 6.2, their common value is equal to the Shannon capacity of the graph. This gives the following closed-form expression for the Shannon capacity of $q$-Kneser graphs.
@@
The Shannon capacity of the $q$-Kneser graph $\operatorname{\mathsf{K}}_{q}(n,k)$, where $n\geq 2k$ and $q$ is a prime power, is given by
@@ 7 A new inequality for the capacity of graphs
The following result provides a relation between the Shannon capacity of any strong product of graphs, and the capacity of the disjoint union of the component graphs. If these component graphs are connected, then their strong product is a connected graph on a number of vertices that is equal to the product of the number of vertices in each component graph, whereas the disjoint union of these component graphs is a disconnected graph on a number of vertices that is equal to the sum of the orders of the component graphs (the latter order is typically much smaller than the former). The motivation for our inequality comes from the following result.
@@
Next, we provide and prove the main result of this section.
@@
Let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{\ell}$ be simple graphs. Then,
@@
In particular, if for every $i\in[\ell]$, one of the following statements holds:
@@
and, thus, by combining (239) and (240), it follows that
@@
Raising both sides of the inequality to the power $\frac{1}{k}$ and letting $k$ tend to infinity gives
@@
An anonymous reviewer brought to our attention that inequality (237) can be related to Eq. (1) in the preprint [47] for the complementary graph entropy, via the results of [48, Lemma 1] and [6, Equation (11.2)].
The next result strengthens [7, Theorem 2.1] by identifying several sufficient conditions under which the inequality holds with equality.
@@
with equality in (245) if the graph $\mathsf{G}$ is either self-complementary and vertex-transitive, self-complementary and strongly regular, a conference graph, a Latin square graph, or the complement of any of these graphs.
@@
Since $\bigl\{(1,1),\ldots,(n,n)\bigr\}$ is a trivial independent set of $\mathsf{G}\boxtimes\overline{\mathsf{G}}$, it follows that
@@
Combining inequalities (246) and (247) yields (245). Furthermore, suppose that the graph $\mathsf{G}$ is one of the following: (1) self-complementary and vertex-transitive, (2) self-complementary and strongly regular, (3) a conference graph, (4) a Latin square graph, or (5) the complement of any of the graphs in (1)–(4). Then, by Theorem 7.2 and [39, Theorems 3.23, 3.26, and 3.28], inequalities (246) and (247) hold with equality, and consequently inequality (245) also holds with equality. ∎
@@
The original proof of inequality (245) in [7, Theorem 2.1] is presented in Appendix E. Our proof of Corollary 7.1 follows a different approach, which enables us to identify sufficient conditions for equality in (245).
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, let $m_{1},\ldots,m_{\ell}\in\mathbb{N}$, and let $m=\sqrt[\ell]{m_{1}\ldots m_{\ell}}$. Then,
@@
By Corollary 7.2 (note that the strong product in the next line involves $k$ graphs),
@@
Calculation of $m$ on the right-hand side of (257) gives, by (253) and (258)–(261),
@@
This is true because if we choose $k$ that is not a multiple of $\mathrm{lcm}(q_{1},\ldots,q_{\ell})$, then there exists $i\in[\ell]$ such that $\alpha_{i}\notin\mathbb{N}$.
@@
$\vartheta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}+\frac{(\vartheta(\mathsf{G})-\sqrt{n})^{2}}{\vartheta(\mathsf{G})},$ (265)
and for every $d$-regular graph $\mathsf{G}$ whose eigenvalues are ordered in non-increasing order as
$\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n},$ (266)
the following holds:
$2+\frac{n-d-1}{1+\lambda_{2}}-\frac{d}{\lambda_{n}}$ $\leq\vartheta(\mathsf{G}+\overline{\mathsf{G}})$
$\leq\frac{n(1+\lambda_{2})}{n-d+\lambda_{2}}-\frac{n\lambda_{n}}{d-\lambda_{n}}.$ (267)
Furthermore,
@@
    If $\mathsf{G}$ is either vertex-transitive or strongly regular, then inequality (265) holds with equality.
@@
    If $\mathsf{G}$ is strongly regular, then the first inequality in (267) is attained with equality.
@@
    If both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then the second inequality in (267) is attained with equality.
@@
$\vartheta(\mathsf{G}+\overline{\mathsf{G}})=\vartheta(\mathsf{G})+\vartheta(\overline{\mathsf{G}})\leq\frac{n(1+\lambda_{2})}{n-d+\lambda_{2}}-\frac{n\lambda_{n}}{d-\lambda_{n}},$ (269)
and
@@
which proves (267). Finally, by [36, Proposition 1], if both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then inequality (269) holds with equality. Moreover, if $\mathsf{G}$ is strongly regular, then inequality (7) also holds with equality. This completes the proof of the stated sufficient conditions for the attainability with equality of the two inequalities in (267). ∎